# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2024-125 LHCb-PAPER-2024-007

### v1

@@ Acknowledgements
We would like to thank Dean Robinson and Michele Papucci for extensive discussions and technical advice on the use of the HAMMER tool. We express our gratitude to our colleagues in the CERN accelerator departments for the excellent performance of the LHC. We thank the technical and administrative staff at the LHCb institutes. We acknowledge support from CERN and from the national agencies: CAPES, CNPq, FAPERJ and FINEP (Brazil); MOST and NSFC (China); CNRS/IN2P3 (France); BMBF, DFG and MPG (Germany); INFN (Italy); NWO (Netherlands); MNiSW and NCN (Poland); MCID/IFA (Romania); MICINN (Spain); SNSF and SER (Switzerland); NASU (Ukraine); STFC (United Kingdom); DOE NP and NSF (USA). We acknowledge the computing resources that are provided by CERN, IN2P3 (France), KIT and DESY (Germany), INFN (Italy), SURF (Netherlands), PIC (Spain), GridPP (United Kingdom), CSCS (Switzerland), IFIN-HH (Romania), CBPF (Brazil), and Polish WLCG (Poland). We are indebted to the communities behind the multiple open-source software packages on which we depend. Individual groups or members have received support from ARC and ARDC (Australia); Key Research Program of Frontier Sciences of CAS, CAS PIFI, CAS CCEPP, Fundamental Research Funds for the Central Universities, and Sci. & Tech. Program of Guangzhou (China); Minciencias (Colombia); EPLANET, Marie Skłodowska-Curie Actions, ERC and NextGenerationEU (European Union); A*MIDEX, ANR, IPhU and Labex P2IO, and Région Auvergne-Rhône-Alpes (France); AvH Foundation (Germany); ICSC (Italy); GVA, XuntaGal, GENCAT, Inditex, InTalent and Prog. Atracción Talento, CM (Spain); SRC (Sweden); the Leverhulme Trust, the Royal Society and UKRI (United Kingdom).
F. Abudinén55
T. Ackernley59
A. A. Adefisoye67
B. Adeva45
M. Adinolfi53
P. Adlarson79
C. Agapopoulou13
C.A. Aidala80
@@
S. Akar64
K. Akiba36
P. Albicocco26
J. Albrecht18
F. Alessio47
M. Alexander58
Z. Aliouche61
P. Alvarez Cartelle54
R. Amalric15
S. Amato3
J.L. Amey53
Y. Amhis13,47
L. An6
L. Anderlini25
M. Andersson49
A. Andreianov42
P. Andreola49
M. Andreotti24
D. Andreou67
A. Anelli29,p
D. Ao7
F. Archilli35,v
M. Argenton24
S. Arguedas Cuendis9
A. Artamonov42
M. Artuso67
E. Aslanides12
M. Atzeni63
B. Audurier14
D. Bacher62
I. Bachiller Perea10
S. Bachmann20
M. Bachmayer48
J.J. Back55
P. Baladron Rodriguez45
V. Balagura14
W. Baldini24
H. Bao7
J. Baptista de Souza Leite59
M. Barbetti25,m
I. R. Barbosa68
R.J. Barlow61
M. Barnyakov23
S. Barsuk13
W. Barter57
M. Bartolini54
J. Bartz67
F. Baryshnikov42
J.M. Basels16
G. Bassi33
B. Batsukh5
A. Bay48
A. Beck55
M. Becker18
F. Bedeschi33
I.B. Bediaga2
S. Belin45
V. Bellee49
K. Belous42
I. Belov27
I. Belyaev34
G. Benane12
G. Bencivenni26
E. Ben-Haim15
A. Berezhnoy42
R. Bernet49
S. Bernet Andres43
A. Bertolin31
C. Betancourt49
F. Betti57
J. Bex54
Ia. Bezshyiko49
J. Bhom39
M.S. Bieker18
N.V. Biesuz24
P. Billoir15
A. Biolchini36
M. Birch60
F.C.R. Bishop10
A. Bitadze61
A. Bizzeti
T. Blake55
F. Blanc48
J.E. Blank18
S. Blusk67
V. Bocharnikov42
J.A. Boelhauve18
O. Boente Garcia14
T. Boettcher64
A. Bohare57
A. Boldyrev42
C.S. Bolognani76
R. Bolzonella24,l
N. Bondar42
F. Borgato31,q,47
S. Borghi61
M. Borsato29,p
J.T. Borsuk39
S.A. Bouchiba48
T.J.V. Bowcock59
A. Boyer47
C. Bozzi24
@@
A. Brea Rodriguez48
N. Breer18
J. Brodzicka39
A. Brossa Gonzalo45
J. Brown59
D. Brundu30
@@ E. Buchanan57,
A. Buonaura49
L. Buonincontri31,q
A.T. Burke61
C. Burr47
A. Butkevich42
J.S. Butter54
J. Buytaert47
W. Byczynski47
S. Cadeddu30
@@ H. Cai72,
R. Calabrese24,l
S. Calderon Ramirez9
L. Calefice44
S. Cali26
M. Calvi29,p
M. Calvo Gomez43
P. Camargo Magalhaes2,z
J. I. Cambon Bouzas45
P. Campana26
D.H. Campora Perez76
A.F. Campoverde Quezada7
S. Capelli29
L. Capriotti24
R. Caravaca-Mora9
A. Carbone23,j
L. Carcedo Salgado45
R. Cardinale27,n
A. Cardini30
P. Carniti29,p
@@ A. Cardini30
A. Casais Vidal63
R. Caspary20
G. Casse59
J. Castro Godinez9
M. Cattaneo47
G. Cavallero24,47
V. Cavallini24,l
S. Celani20
D. Cervenkov62
S. Cesare28,o
A.J. Chadwick59
I. Chahrour80
M. Charles15
Ph. Charpentier47
C.A. Chavez Barajas59
M. Chefdeville10
C. Chen12
S. Chen5
Z. Chen7
A. Chernov39
S. Chernyshenko51
V. Chobanova78
S. Cholak48
M. Chrzaszcz39
A. Chubykin42
V. Chulikov42
P. Ciambrone26
X. Cid Vidal45
G. Ciezarek47
P. Cifra47
P.E.L. Clarke57
M. Clemencic47
H.V. Cliff54
J. Closier47
C. Cocha Toapaxi20
V. Coco47
J. Cogan12
E. Cogneras11
L. Cojocariu41
P. Collins47
T. Colombo47
A. Comerma-Montells44
L. Congedo22
A. Contu30
N. Cooke58
I. Corredoira 45
A. Correia15
G. Corti47
@@
B. Couturier47
D.C. Craik49
M. Cruz Torres2,g
E. Curras Rivera48
R. Currie57
C.L. Da Silva66
S. Dadabaev42
L. Dai69
X. Dai6
E. Dall’Occo18
J. Dalseno45
C. D’Ambrosio47
J. Daniel11
A. Danilina42
P. d’Argent22
A. Davidson55
J.E. Davies61
A. Davis61
O. De Aguiar Francisco61
C. De Angelis30,k
F. De Benedetti47
J. de Boer36
K. De Bruyn75
S. De Capua61
M. De Cian20,47
U. De Freitas Carneiro Da Graca2,b
E. De Lucia26
J.M. De Miranda2
L. De Paula3
M. De Serio22,h
P. De Simone26
F. De Vellis18
J.A. de Vries76
F. Debernardis22
D. Decamp10
V. Dedu12
L. Del Buono15
B. Delaney63
H.-P. Dembinski18
J. Deng8
V. Denysenko49
O. Deschamps11
F. Dettori30,k
B. Dey74
P. Di Nezza26
I. Diachkov42
S. Didenko42
S. Ding67
L. Dittmann20
V. Dobishuk51
A. D. Docheva58
C. Dong4
A.M. Donohoe21
F. Dordei30
A.C. dos Reis2
A. D. Dowling67
W. Duan70
P. Duda77
M.W. Dudek39
L. Dufour47
V. Duk32
P. Durante47
M. M. Duras77
J.M. Durham66
O. D. Durmus74
A. Dziurda39
A. Dzyuba42
S. Easo56
@@ E. Eckstein17,
U. Egede1
A. Egorychev42
V. Egorychev42
S. Eisenhardt57
E. Ejopu61
L. Eklund79
M. Elashri64
J. Ellbracht18
S. Ely60
A. Ene41
E. Epple64
J. Eschle67
S. Esen20
T. Evans61
F. Fabiano30,k,47
L.N. Falcao2
Y. Fan7
B. Fang72
L. Fantini32,r
M. Faria48
K. Farmer57
D. Fazzini29,p
L. Felkowski77
M. Feng5,7
M. Feo18,47
M. Fernandez Gomez45
A.D. Fernez65
F. Ferrari23
F. Ferreira Rodrigues3
M. Ferrillo49
M. Ferro-Luzzi47
S. Filippov42
R.A. Fini22
M. Fiorini24,l
K.M. Fischer62
D.S. Fitzgerald80
C. Fitzpatrick61
F. Fleuret14
M. Fontana23
L. F. Foreman61
R. Forty47
D. Foulds-Holt54
M. Franco Sevilla65
M. Frank47
E. Franzoso24,l
G. Frau20
C. Frei47
D.A. Friday61
J. Fu7
Q. Fuehring18
Y. Fujii1
T. Fulghesu15
E. Gabriel36
G. Galati22
M.D. Galati36
A. Gallas Torreira45
D. Galli23,j
S. Gambetta57
M. Gandelman3
P. Gandini28
B. Ganie61
H. Gao7
R. Gao62
Y. Gao8
Y. Gao6
@@ H. Gao7
M. Garau30,k
L.M. Garcia Martin48
P. Garcia Moreno44
J. García Pardiñas47
K. G. Garg8
L. Garrido44
C. Gaspar47
R.E. Geertsema36
L.L. Gerken18
E. Gersabeck61
M. Gersabeck61
T. Gershon55
@@ E. Gersabeck61
L. Giambastiani31,q
F. I. Giasemis15,e
V. Gibson54
H.K. Giemza40
A.L. Gilman62
M. Giovannetti26
A. Gioventù44
P. Gironella Gironell44
C. Giugliano24,l
M.A. Giza39
E.L. Gkougkousis60
F.C. Glaser13,20
V.V. Gligorov15,47
C. Göbel68
E. Golobardes43
D. Golubkov42
A. Golutvin60,42,47
A. Gomes2,a,†
S. Gomez Fernandez44
F. Goncalves Abrantes62
M. Goncerz39
G. Gong4
J. A. Gooding18
I.V. Gorelov42
C. Gotti29
J.P. Grabowski17
L.A. Granado Cardoso47
E. Graugés44
E. Graverini48,t
L. Grazette55
G. Graziani
A. T. Grecu41
L.M. Greeven36
N.A. Grieser64
L. Grillo58
S. Gromov42
C. Gu14
M. Guarise24
M. Guittiere13
V. Guliaeva42
P. A. Günther20
A.-K. Guseinov48
E. Gushchin42
Y. Guz6,42,47
T. Gys47
K. Habermann17
T. Hadavizadeh1
C. Hadjivasiliou65
G. Haefeli48
C. Haen47
J. Haimberger47
@@ C. Haen47
M.M. Halvorsen47
P.M. Hamilton65
J. Hammerich59
Q. Han8
X. Han20
S. Hansmann-Menzemer20
L. Hao7
N. Harnew62
M. Hartmann13
J. He7,c
F. Hemmer47
C. Henderson64
R.D.L. Henderson1,55
A.M. Hennequin47
K. Hennessy59
L. Henry48
J. Herd60
P. Herrero Gascon20
J. Heuel16
A. Hicheur3
@@ G. Hijano Mendizabal49,
D. Hill48
S.E. Hollitt18
J. Horswill61
R. Hou8
Y. Hou11
@@ D. Hill48
J. Hu70
W. Hu6
X. Hu4
W. Huang7
W. Hulsbergen36
R.J. Hunter55
M. Hushchyn42
D. Hutchcroft59
D. Ilin42
P. Ilten64
A. Inglessi42
A. Iniukhin42
A. Ishteev42
K. Ivshin42
R. Jacobsson47
H. Jage16
S.J. Jaimes Elles46,73
S. Jakobsen47
E. Jans36
B.K. Jashal46
A. Jawahery65,47
V. Jevtic18
E. Jiang65
X. Jiang5,7
Y. Jiang7
Y. J. Jiang6
M. John62
D. Johnson52
C.R. Jones54
T.P. Jones55
S. Joshi40
B. Jost47
N. Jurik47
I. Juszczak39
D. Kaminaris48
S. Kandybei50
Y. Kang4
C. Kar11
M. Karacson47
D. Karpenkov42
A. Kauniskangas48
J.W. Kautz64
F. Keizer47
M. Kenzie54
T. Ketel36
B. Khanji67
A. Kharisova42
S. Kholodenko33,47
G. Khreich13
T. Kirn16
V.S. Kirsebom29,p
O. Kitouni63
S. Klaver37
N. Kleijne33,s
K. Klimaszewski40
M.R. Kmiec40
S. Koliiev51
L. Kolk18
A. Konoplyannikov42
P. Kopciewicz38,47
P. Koppenburg36
M. Korolev42
I. Kostiuk36

### v2

@@ Acknowledgements
We would like to thank Dean Robinson and Michele Papucci for extensive discussions and technical advice on the use of the HAMMER tool. We express our gratitude to our colleagues in the CERN accelerator departments for the excellent performance of the LHC. We thank the technical and administrative staff at the LHCb institutes. We acknowledge support from CERN and from the national agencies: CAPES, CNPq, FAPERJ and FINEP (Brazil); MOST and NSFC (China); CNRS/IN2P3 (France); BMBF, DFG and MPG (Germany); INFN (Italy); NWO (Netherlands); MNiSW and NCN (Poland); MCID/IFA (Romania); MICINN (Spain); SNSF and SER (Switzerland); NASU (Ukraine); STFC (United Kingdom); DOE NP and NSF (USA). We acknowledge the computing resources that are provided by CERN, IN2P3 (France), KIT and DESY (Germany), INFN (Italy), SURF (Netherlands), PIC (Spain), GridPP (United Kingdom), CSCS (Switzerland), IFIN-HH (Romania), CBPF (Brazil), and Polish WLCG (Poland). We are indebted to the communities behind the multiple open-source software packages on which we depend. Individual groups or members have received support from ARC and ARDC (Australia); Key Research Program of Frontier Sciences of CAS, CAS PIFI, CAS CCEPP, Fundamental Research Funds for the Central Universities, and Sci. & Tech. Program of Guangzhou (China); Minciencias (Colombia); EPLANET, Marie SkĹodowska-Curie Actions, ERC and NextGenerationEU (European Union); A*MIDEX, ANR, IPhU and Labex P2IO, and RĂŠgion Auvergne-RhĂ´ne-Alpes (France); AvH Foundation (Germany); ICSC (Italy); GVA, XuntaGal, GENCAT, Inditex, InTalent and Prog. AtracciĂłn Talento, CM (Spain); SRC (Sweden); the Leverhulme Trust, the Royal Society and UKRI (United Kingdom).
F. AbudinĂŠn55â
T. Ackernley59â
A. A. Â Adefisoye67â
B. Adeva45â
M. Adinolfi53â
P. Adlarson79â
C. Agapopoulou13â
C.A. Aidala80â
@@
S. Akar64â
K. Akiba36â
P. Albicocco26â
J. Albrecht18â
F. Alessio47â
M. Alexander58â
Z. Aliouche61â
P. Alvarez Cartelle54â
R. Amalric15â
S. Amato3â
J.L. Amey53â
Y. Amhis13,47â
L. An6â
L. Anderlini25â
M. Andersson49â
A. Andreianov42â
P. Andreola49â
M. Andreotti24â
D. Andreou67â
A. Anelli29,pâ
D. Ao7â
F. Archilli35,vâ
M. Argenton24â
S. Arguedas Cuendis9â
A. Artamonov42â
M. Artuso67â
E. Aslanides12â
M. Atzeni63â
B. Audurier14â
D. Bacher62â
I. Bachiller Perea10â
S. Bachmann20â
M. Bachmayer48â
J.J. Back55â
P. Baladron Rodriguez45â
V. Balagura14â
W. Baldini24â
H. Â Bao7â
J. Baptista de Souza Leite59â
M. Barbetti25,mâ
I. R. Barbosa68â
R.J. Barlow61â
M. Barnyakov23â
S. Barsuk13â
W. Barter57â
M. Bartolini54â
J. Bartz67â
F. Baryshnikov42â
J.M. Basels16â
G. Bassi33â
B. Batsukh5â
A. Bay48â
A. Beck55â
M. Becker18â
F. Bedeschi33â
I.B. Bediaga2â
S. Belin45â
V. Bellee49â
K. Belous42â
I. Belov27â
I. Belyaev34â
G. Benane12â
G. Bencivenni26â
E. Ben-Haim15â
A. Berezhnoy42â
R. Bernet49â
S. Bernet Andres43â
A. Bertolin31â
C. Betancourt49â
F. Betti57â
J. Â Bex54â
Ia. Bezshyiko49â
J. Bhom39â
M.S. Bieker18â
N.V. Biesuz24â
P. Billoir15â
A. Biolchini36â
M. Birch60â
F.C.R. Bishop10â
A. Bitadze61â
A. Bizzetiâ
T. Blake55â
F. Blanc48â
J.E. Blank18â
S. Blusk67â
V. Bocharnikov42â
J.A. Boelhauve18â
O. Boente Garcia14â
T. Boettcher64â
A. Â Bohare57â
A. Boldyrev42â
C.S. Bolognani76â
R. Bolzonella24,lâ
N. Bondar42â
F. Borgato31,q,47â
S. Borghi61â
M. Borsato29,pâ
J.T. Borsuk39â
S.A. Bouchiba48â
T.J.V. Bowcock59â
A. Boyer47â
C. Bozzi24â
@@
A. Brea Rodriguez48â
N. Breer18â
J. Brodzicka39â
A. Brossa Gonzalo45â
J. Brown59â
D. Brundu30â
@@ E. Buchanan57,
A. Buonaura49â
L. Buonincontri31,qâ
A.T. Burke61â
C. Burr47â
A. Butkevich42â
J.S. Butter54â
J. Buytaert47â
W. Byczynski47â
S. Cadeddu30â
@@ H. Cai72,
R. Calabrese24,lâ
S. Calderon Ramirez9â
L. Calefice44â
S. Cali26â
M. Calvi29,pâ
M. Calvo Gomez43â
P. Camargo Magalhaes2,zâ
J. I. Cambon Bouzas45â
P. Campana26â
D.H. Campora Perez76â
A.F. Campoverde Quezada7â
S. Capelli29â
L. Capriotti24â
R. Caravaca-Mora9â
A. Carbone23,jâ
L. Carcedo Salgado45â
R. Cardinale27,nâ
A. Cardini30â
P. Carniti29,pâ
@@ A. Cardini30â
A. Casais Vidal63â
R. Caspary20â
G. Casse59â
J. Castro Godinez9â
M. Cattaneo47â
G. Cavallero24,47â
V. Cavallini24,lâ
S. Celani20â
D. Cervenkov62â
S. Â Cesare28,oâ
A.J. Chadwick59â
I. Chahrour80â
M. Charles15â
Ph. Charpentier47â
C.A. Chavez Barajas59â
M. Chefdeville10â
C. Chen12â
S. Chen5â
Z. Chen7â
A. Chernov39â
S. Chernyshenko51â
V. Chobanova78â
S. Cholak48â
M. Chrzaszcz39â
A. Chubykin42â
V. Chulikov42â
P. Ciambrone26â
X. Cid Vidal45â
G. Ciezarek47â
P. Cifra47â
P.E.L. Clarke57â
M. Clemencic47â
H.V. Cliff54â
J. Closier47â
C. Cocha Toapaxi20â
V. Coco47â
J. Cogan12â
E. Cogneras11â
L. Cojocariu41â
P. Collins47â
T. Colombo47â
A. Comerma-Montells44â
L. Congedo22â
A. Contu30â
N. Cooke58â
I. Corredoira 45â
A. Correia15â
G. Corti47â
@@
B. Couturier47â
D.C. Craik49â
M. Cruz Torres2,gâ
E. Curras Rivera48â
R. Currie57â
C.L. Da Silva66â
S. Dadabaev42â
L. Dai69â
X. Dai6â
E. DallâOcco18â
J. Dalseno45â
C. DâAmbrosio47â
J. Daniel11â
A. Danilina42â
P. dâArgent22â
A. Â Davidson55â
J.E. Davies61â
A. Davis61â
O. De Aguiar Francisco61â
C. De Angelis30,kâ
F. De Benedetti47â
J. de Boer36â
K. De Bruyn75â
S. De Capua61â
M. De Cian20,47â
U. De Freitas Carneiro Da Graca2,bâ
E. De Lucia26â
J.M. De Miranda2â
L. De Paula3â
M. De Serio22,hâ
P. De Simone26â
F. De Vellis18â
J.A. de Vries76â
F. Debernardis22â
D. Decamp10â
V. Dedu12â
L. Del Buono15â
B. Delaney63â
H.-P. Dembinski18â
J. Deng8â
V. Denysenko49â
O. Deschamps11â
F. Dettori30,kâ
B. Dey74â
P. Di Nezza26â
I. Diachkov42â
S. Didenko42â
S. Ding67â
L. Dittmann20â
V. Dobishuk51â
A. D. Â Docheva58â
C. Dong4â
A.M. Donohoe21â
F. Dordei30â
A.C. dos Reis2â
A. D. Â Dowling67â
W. Duan70â
P. Duda77â
M.W. Dudek39â
L. Dufour47â
V. Duk32â
P. Durante47â
M. M. Duras77â
J.M. Durham66â
O. D. Â Durmus74â
A. Dziurda39â
A. Dzyuba42â
S. Easo56â
@@ E. Eckstein17,
U. Egede1â
A. Egorychev42â
V. Egorychev42â
S. Eisenhardt57â
E. Ejopu61â
L. Eklund79â
M. Elashri64â
J. Ellbracht18â
S. Ely60â
A. Ene41â
E. Epple64â
J. Eschle67â
S. Esen20â
T. Evans61â
F. Fabiano30,k,47â
L.N. Falcao2â
Y. Fan7â
B. Fang72â
L. Fantini32,râ
M. Faria48â
K. Â Farmer57â
D. Fazzini29,pâ
L. Felkowski77â
M. Feng5,7â
M. Feo18,47â
M. Fernandez Gomez45â
A.D. Fernez65â
F. Ferrari23â
F. Ferreira Rodrigues3â
M. Ferrillo49â
M. Ferro-Luzzi47â
S. Filippov42â
R.A. Fini22â
M. Fiorini24,lâ
K.M. Fischer62â
D.S. Fitzgerald80â
C. Fitzpatrick61â
F. Fleuret14â
M. Fontana23â
L. F. Â Foreman61â
R. Forty47â
D. Foulds-Holt54â
M. Franco Sevilla65â
M. Frank47â
E. Franzoso24,lâ
G. Frau20â
C. Frei47â
D.A. Friday61â
J. Fu7â
Q. Fuehring18â
Y. Fujii1â
T. Fulghesu15â
E. Gabriel36â
G. Galati22â
M.D. Galati36â
A. Gallas Torreira45â
D. Galli23,jâ
S. Gambetta57â
M. Gandelman3â
P. Gandini28â
B. Â Ganie61â
H. Gao7â
R. Gao62â
Y. Gao8â
Y. Gao6â
@@ H. Gao7â
M. Garau30,kâ
L.M. Garcia Martin48â
P. Garcia Moreno44â
J. GarcĂa PardiĂąas47â
K. G. Â Garg8â
L. Garrido44â
C. Gaspar47â
R.E. Geertsema36â
L.L. Gerken18â
E. Gersabeck61â
M. Gersabeck61â
T. Gershon55â
@@ E. Gersabeck61â
L. Giambastiani31,qâ
F. I. Giasemis15,eâ
V. Gibson54â
H.K. Giemza40â
A.L. Gilman62â
M. Giovannetti26â
A. GioventĂš44â
P. Gironella Gironell44â
C. Giugliano24,lâ
M.A. Giza39â
E.L. Gkougkousis60â
F.C. Glaser13,20â
V.V. Gligorov15,47â
C. GĂśbel68â
E. Golobardes43â
D. Golubkov42â
A. Golutvin60,42,47â
A. Gomes2,a,â â
S. Gomez Fernandez44â
F. Goncalves Abrantes62â
M. Goncerz39â
G. Gong4â
J. A. Gooding18â
I.V. Gorelov42â
C. Gotti29â
J.P. Grabowski17â
L.A. Granado Cardoso47â
E. GraugĂŠs44â
E. Graverini48,tâ
L. Grazette55â
G. Grazianiâ
A. T. Grecu41â
L.M. Greeven36â
N.A. Grieser64â
L. Grillo58â
S. Gromov42â
C. Â Gu14â
M. Guarise24â
M. Guittiere13â
V. Guliaeva42â
P. A. GĂźnther20â
A.-K. Guseinov48â
E. Gushchin42â
Y. Guz6,42,47â
T. Gys47â
K. Habermann17â
T. Hadavizadeh1â
C. Hadjivasiliou65â
G. Haefeli48â
C. Haen47â
J. Haimberger47â
@@ C. Haen47â
M.M. Halvorsen47â
P.M. Hamilton65â
J. Hammerich59â
Q. Han8â
X. Han20â
S. Hansmann-Menzemer20â
L. Hao7â
N. Harnew62â
M. Hartmann13â
J. He7,câ
F. Hemmer47â
C. Henderson64â
R.D.L. Henderson1,55â
A.M. Hennequin47â
K. Hennessy59â
L. Henry48â
J. Herd60â
P. Herrero Gascon20â
J. Heuel16â
A. Hicheur3â
@@ G. Hijano Mendizabal49,
D. Hill48â
S.E. Hollitt18â
J. Horswill61â
R. Hou8â
Y. Hou11â
@@ D. Hill48â
J. Hu70â
W. Hu6â
X. Hu4â
W. Huang7â
W. Hulsbergen36â
R.J. Hunter55â
M. Hushchyn42â
D. Hutchcroft59â
D. Ilin42â
P. Ilten64â
A. Inglessi42â
A. Iniukhin42â
A. Ishteev42â
K. Ivshin42â
R. Jacobsson47â
H. Jage16â
S.J. Jaimes Elles46,73â
S. Jakobsen47â
E. Jans36â
B.K. Jashal46â
A. Jawahery65,47â
V. Jevtic18â
E. Jiang65â
X. Jiang5,7â
Y. Jiang7â
Y. J. Â Jiang6â
M. John62â
D. Johnson52â
C.R. Jones54â
T.P. Jones55â
S. Joshi40â
B. Jost47â
N. Jurik47â
I. Juszczak39â
D. Kaminaris48â
S. Kandybei50â
Y. Kang4â
C. Kar11â
M. Karacson47â
D. Karpenkov42â
A. Kauniskangas48â
J.W. Kautz64â
F. Keizer47â
M. Kenzie54â
T. Ketel36â
B. Khanji67â
A. Kharisova42â
S. Kholodenko33,47â
G. Khreich13â
T. Kirn16â
V.S. Kirsebom29,pâ
O. Kitouni63â
S. Klaver37â
N. Kleijne33,sâ
K. Klimaszewski40â
M.R. Kmiec40â
S. Koliiev51â
L. Kolk18â
A. Konoplyannikov42â
P. Kopciewicz38,47â
P. Koppenburg36â
M. Korolev42â
I. Kostiuk36â